Consider the titration of 50.00 mL of a mixture of NaI and NaBr with 0.0500 M silver nitrate. The first equivalence point was observed at 10.50 mL and the second at 25.50 mL. Calculate the molar concentration of the two salts in the sample


1
Expert's answer
2021-07-02T07:26:01-0400

Moles of mixture = "\frac{50.00}{1000}×0.0500=2.5×10^{-3}" Moles



Molarity = "\frac{Moles }{Volume}"


Molarity = "\frac{2.5×10^{-3}}{0.0105}=0.238M"


Molarity "=\frac{2.5×10^{-3}}{0.0255}=0.098M"